# Compressible and immiscible fluids with arbitrary density ratios: diffuse interface theory

### v2

@@ I Introduction
In this work, we explore the high-density ratio problem through the physical principle of energy minimization. We propose a generalized theoretical framework for the evolution equations of fluid dynamics and derive a novel kinetic equation governing density dynamics at the water-air interface. The resulting density evolution equation yields three key contributions: (A): A universal formulation for sound speed; (B): A generalized equation of state for an ideal gas; (C): A generalization of Bernoulli’s law, extending the conservation of potential and kinetic energy. For each contribution, a case study and application is presented in Section VII.
The rest of the paper is structured as follows. In Section II, we provide a brief overview of two high-density ratio models from the literature, clarify the distinction between the total time derivative and the mass conservative derivative, and examine the validity of the conventional continuity equation. Section III outlines the setup of our study and introduces a new concept of volume concentration. In Section IV, we present a dissipation-conservation principle that establishes a connection between mass conservation and energy dissipation. Section V applies this principle to derive the classical fluid dynamics equations for uniform systems. Building on these results, Section VI extends the derivation to non-uniform systems, derives the energy dissipation, and formulates a novel kinetic equation for density. In Section VII, we demonstrate the application of our theory by deriving the sound speed, the equation of state (EOS) for an ideal gas, and the density profile across a water-air interface. Finally, Section VIII concludes the paper with a summary of our findings.
@@ II.2 Abels-Garcke-Grün approach
Other existing literatures are based either on the continuity equation for the density evolution Tryggvason et al. (2001); Puckett et al. (1997); Fedkiw et al. (1999) (which will be discussed in Section II C) or on the linear interpolation of the density Fakhari et al. (2017); Shukla et al. (2010) according to Eq. (3).
@@ III Setup, definition and excess volume concept
where $v_{e}\neq 0$ denotes the excess volume, which exhibits spatial variation as a result of local pressure changes. Instead of explicitly solving for $v_{e}$, we redistribute it between $v_{w}$ and $v_{a}$ by an appropriate allocation scheme, resulting in two subregions with volumes $v_{w}^{\prime}$ and $v_{a}^{\prime}$. Consequently, the total volume satisfies $V=v_{w}^{\prime}+v_{a}^{\prime}$ (Fig. 2b). This redistribution of excess volume implies that the local partial densities of water and gas must vary accordingly, since the individual masses of the two components within the volume element $V$ remain constant. We define the volume composition as
@@
The partial densities $\rho_{w}^{\prime}$ and $\rho_{a}^{\prime}$ are implicitly influenced by the excess volume and are therefore no longer constant, as they vary with the redistribution of the excess volume. Consequently, to fully characterize the state of the system, two evolution equations are required: one governing the volume composition $\phi$, and the other governing the density $\rho$.
The classical treatment of incompressible flow corresponds to a special case in which the excess volume vanishes both in the bulk phases and across the interface. Under this assumption, it is sufficient to describe the system using a single variable—either the density $\rho$ or the volume composition $\phi$. A representative example employing $\phi$ is the van der Waals formulation of free energy and its associated diffuse-interface framework Wöhrwag et al. (2018); Ashour et al. (2023); Bestehorn et al. (2021); Jacqmin (1999); Anderson et al. (1998); Yue et al. (2004); Lowengrub and Truskinovsky (1998), in which a single phase order parameter $\phi$ is used to characterize the system’s state and evolution. This well-established approach is commonly referred to as the phase-field model or Cahn–Hilliard model. Another classical example is Euler’s formulation for density evolution, where the standard continuity equation, Eq. (26), is adopted. For incompressible flow, Euler’s formulation reduces to the condition that the convective velocity field is divergence-free. It is important to note that, for incompressible fluids, the van der Waals and Euler formulations cannot be applied simultaneously, as doing so would lead to an overdetermined system.
@@ IV Dissipation principle
where $\tau_{0}$ is a positive constant. Substituting Eq. (34) into Eq. (33), integrating by parts, and using the no-flux boundary condition, we obtain
@@
Thus, the coefficient $\tau_{0}$ establishes a connection between the purely dynamic parameter $\mathcal{D}$ (independent of conservative energy) and the energy-dependent term $\frac{\mathrm{d}^{2}e}{\mathrm{d}u^{2}}$ via the relation $\tau_{0}=\mathcal{D}/\frac{\mathrm{d}^{2}e}{\mathrm{d}u^{2}}$. In other words, $\tau_{0}$ represents the ratio of non-conservative to conservative energy contributions to the kinetics of the system.
@@ V Uniform system: Derivation of Euler’s equation and Newton’s law
The free energy density depends on the volume concentration $\phi$ and temperature $T$, and may be expressed as $f=\theta-Ts$, where $\theta$ denotes the internal energy and $s\sim\phi\ln\phi$ represents the mixing entropy, which is distinct from the thermal (heat) entropy.
@@
where the mass diffusivity $D$ is defined in terms of the mobility $M$ as $D=M(\mathrm{d}^{2}f/\mathrm{d}\phi^{2})$.
@@ VI.1 Energy formulation and dissipation
In contrast to the uniform system, the free energy of a non-uniform system depends not only on $\phi$ but also on $\nabla\phi$, as derived by Cahn Cahn and Hilliard (1958) and van der Waals Rowlinson (1979), known as van der Waals-Cahn theory. The integration over the local free energy density $e(\bm{\phi},\nabla\bm{\phi})$ leads to
$\mathcal{F}(\bm{\phi},\nabla\bm{\phi})=\int_{\Omega}e(\bm{\phi},\nabla\bm{\phi})d\Omega,$ (53)
$e(\bm{\phi},\nabla\bm{\phi})=f(\bm{\phi})+\sum_{i=1}^{K}\frac{1}{2}\kappa_{i}(\nabla\phi_{i})^{2},$ (54)
where $f(\bm{\phi})$ depicts the bulk free energy density and $\kappa_{i}$ represents the gradient energy coefficient which is related to the interfacial tension of the $\alpha$-$\beta$ interface, $\sigma_{\alpha\beta}$.
Figure 3: Origin of the pressure energy for uniform and nonuniform systems. For a uniform system (a), the pressure at any position is the same, which is taken as a reference value. In a non-uniform system with density variation (b), the pressure is generally non-uniform due to the asymmetric collisions of different sized molecules. Blue and red spheres depict two molecules or atoms with distinct sizes, leading to a variation of the density in space. The arrows sketch the thermal motion of the molecules/atoms.
The $\mathcal{P}=\int_{\Omega}pd\Omega$ term, often overlooked previously and introduced here, describes the pressure energy. We formulate the pressure energy density as Chen and Doolen (1998)
$-p=e-\sum_{i=1}^{K}\mu_{i}\phi_{i},$ (55)
where $\mu_{i}$ is the chemical potential of species $i$, the formulation of which will be derived in the following via the energy dissipation principle. Herein, an essential consideration is that the pressure $p$ is always associated with the free energy $\mathcal{F}$. A physical interpretation of the pressure is shown in Fig. 3. A homogeneous mixture of two distinct species (Fig. 3a) leads to a uniform pressure $p_{0}(\bm{r})=\text{const},\ \forall\bm{r}\in\Omega$, which can be taken as a reference value. However, when the two species are not homogeneously mixed (Fig. 3b), the intrinsic pressure is non-uniform in space; the non-uniform pressure is caused, for example, by the collisions of different molecules with distinct sizes due to the asymmetry in space. In the special case of a homogeneous mixture, the pressure defined in Eq. (55) is consistent with the Landau potential (sometimes called the thermodynamic pressure Wang et al. (2023)), given by
$\Psi=f(\bm{\phi})-\sum_{i=1}^{K}(\partial f/\partial\phi_{i})\phi_{i}.$ (56)
@@
$\mathcal{K}=\int_{\Omega}\bm{\zeta}\cdot\mathbf{u}d\Omega,$ (57)
@@
The $\mathcal{W}$ term in Eq. (52) represents the wall free energy, which is integrated over the wall free energy density $f_{w}$ on the boundary $\Gamma$ as $\mathcal{W}=\int_{\Gamma}f_{w}d\Gamma$. The formulation of the wall free energy density $f_{w}$ has been comprehensively discussed elsewhere Wang and Nestler (2024); Wang et al. (2023); Wang and Nestler (2021) and is not the scope of this work.
@@
$\dot{\mathcal{F}}(\bm{\phi},\nabla\bm{\phi})=\int_{\Omega}\sum_{i=1}^{K}\frac{\partial e}{\partial\phi_{i}}\frac{\mathrm{d}\phi_{i}}{\mathrm{d}t}+\frac{\partial e}{\partial\nabla\phi_{i}}\cdot\frac{\mathrm{d}\nabla\phi_{i}}{\mathrm{d}t}d\Omega.$ (58)
By using the vector calculus $\nabla\cdot(\nabla\phi_{i}\otimes\nabla\phi_{i})=\nabla^{2}\phi_{i}\nabla\phi_{i}+(\nabla\phi_{i}\cdot\nabla)\nabla\phi_{i}$, the definition of materials derivative $\dot{\nabla\phi_{i}}=\partial_{t}\nabla\phi_{i}+\mathbf{u}\cdot\nabla\nabla\phi_{i}$, and integration by parts with no-flux boundary condition for a closed system (more details can be found in Wang et al. (2023); Zhang et al. (2024)), we obtain the following equations
$\dot{\mathcal{F}}=\int_{\Omega}\sum_{i=1}^{K}\mu_{i}\dot{\phi}_{i}+(\nabla\cdot\underline{\underline{\bm{\Theta}}})\cdot\mathbf{u}d\Omega;$ (59)
$\mu_{i}=\frac{\partial e}{\partial\phi_{i}}-\nabla\cdot\frac{\partial e}{\partial\nabla\phi_{i}};$ (60)
@@
The first part of Eq. (59) defines a generalized formulation for the chemical potential, as stated in Eq. (60). At thermodynamic equilibrium, the chemical potential is a constant value throughout the system. In a special case of a zero chemical potential, i.e., $\mu_{i}=0$, Eq. (60) is equivalent to the Euler-Lagrange equation and replicates the Young-Laplace equation; Eq. (60) also provides a physical interpretation for the variational derivative in mathematics. The second part of Eq. (59) defines a stress tensor $\underline{\underline{\bm{\Theta}}}$, which is responsible for the transformation of the potential free energy into the macroscopic kinetic energy Jacqmin (1999). This stress tensor is consistent with the Korteweg stress by evaluating the derivative via considering Eq. (54), namely $\partial_{\nabla\phi_{i}}e=\kappa_{i}\nabla\phi_{i}$.
@@
where $\mathbf{I}$ is an identity tensor. For the derivation of Eq. (62) and Eq. (63), we have used the chain rule and the definition of the velocity, $\dot{\mathbf{r}}=\mathbf{u}$. Summarizing Eq. (59), Eq. (60), Eq. (61), Eq. (62), and Eq. (63) and collecting all the $\cdot\mathbf{u}$ terms, we obtain the total time derivative of the system energy $\mathcal{E}$ as
$\dot{\mathcal{E}}=\int_{\Omega}\sum_{i=1}^{K}\underbrace{\mu_{i}\dot{\phi_{i}}}_{\text{I}}+\underbrace{[\nabla\cdot(\underline{\underline{\bm{\Theta}}}+p\mathbf{I})+\dot{\bm{\zeta}}]}_{\text{II}}\cdot\mathbf{u}+\underbrace{\bm{\zeta}\cdot\dot{\mathbf{u}}}_{\text{III}}d\Omega.$ (64)
The physical interpretation for the three parts of the energy evolution in Eq. (64) is sketched in Fig. 4.
Figure 4: The energy dissipation mechanisms in Eq. (64). I: The diffusion via the exchange of the molecules and vacancies leads to the dissipation of the free energy $\mathcal{F}$. II: The gradient of the free energy $\mathcal{F}$ and the associated pressure $\mathcal{P}$ result in an increase in the momentum $\bm{\zeta}=\rho\mathbf{u}$. III: The difference in the velocity, density, and viscosity in distinct subdomains (dashed rectangle) gives rise to the dissipation of the velocity $\mathbf{u}$ caused by the ‘spring’ drag effect; the spring symbol denote the interaction between the atoms/molecules in different subdomains.

### v3

@@ I Introduction
In this work, we explore the high-density ratio problem through the physical principle of energy minimization. We propose a generalized theoretical framework for the evolution equations of fluid dynamics and derive a novel kinetic equation governing density dynamics at the fluid-fluid interface. The resulting density evolution equation yields three key contributions: (A): A universal formulation for sound speed; (B): A generalized equation of state for fluids; (C): A generalization of Bernoulli’s law, extending the conservation of potential and kinetic energy. For each contribution, a case study and application is presented in Section VII.
The rest of the paper is structured as follows. In Section II, we provide a brief overview of two high-density ratio models from literature, clarify the distinction between the total time derivative and the mass conservative derivative, and examine the validity of the conventional continuity equation. Section III outlines the setup of our study and introduces a new concept of volume concentration. In Section IV, we present a dissipation-conservation principle that establishes a connection between mass conservation and energy dissipation. Section V applies this principle to derive the classical fluid dynamics equations for uniform systems. Building on these results, Section VI extends the derivation to non-uniform systems, derives the free energy dissipation, and formulates a novel kinetic equation for density. In Section VII, we demonstrate the application of our theory by deriving the sound speed, the equation of state (EOS) for fluids, and the density profile across a water-air interface. Finally, Section VIII concludes the paper with a summary of our findings.
@@ II.2 Abels-Garcke-Grün approach
Other existing literatures are based either on the continuity equation for the density evolution Tryggvason et al. (2001); Puckett et al. (1997); Fedkiw et al. (1999) (which will be discussed in Section II C and Section VI D) or on the linear interpolation of the density Fakhari et al. (2017); Shukla et al. (2010) according to Eq. (3).
@@ III Setup, definition and excess volume concept
where $v_{e}\neq 0$ denotes the excess volume, which exhibits spatial variation as a result of local pressure changes. Instead of explicitly solving for $v_{e}$, we redistribute it between $v_{w}$ and $v_{a}$ by an appropriate allocation scheme, resulting in two subregions with volumes $v_{w}^{\prime}$ and $v_{a}^{\prime}$. Consequently, the total volume satisfies $V=v_{w}^{\prime}+v_{a}^{\prime}$ (Fig. 2b). This redistribution of excess volume implies that the local partial densities of water and gas must vary accordingly, since the individual masses of the two components within the volume element $V$ remain constant. We define the volume composition including the information of excess volume as
@@
Note that both $\rho$ and $\phi$ encode the information associated with the excess volume. The partial densities $\rho_{w}^{\prime}$ and $\rho_{a}^{\prime}$ are implicitly influenced by the excess volume and are therefore no longer constant, as they vary with the redistribution of the excess volume. Consequently, to fully characterize the state of the system, two evolution equations are required: one governing the volume composition $\phi$, and the other governing the density $\rho$.
The classical treatment of incompressible flow corresponds to a special case in which the excess volume vanishes both in the bulk phases and across the interface. Under this assumption, it is sufficient to describe the system using a single variable—either the density $\rho$ or the volume composition $\phi$. A representative example employing $\phi$ is the van der Waals formulation of free energy and its associated diffuse-interface framework Wöhrwag et al. (2018); Ashour et al. (2023); Bestehorn et al. (2021); Jacqmin (1999); Anderson et al. (1998); Yue et al. (2004); Lowengrub and Truskinovsky (1998), in which a single phase order parameter $\phi$ is used to characterize the system’s state and evolution. This well-established approach Ridl and Wagner (2018) is commonly referred to as the phase-field model or Cahn–Hilliard model. Another classical example is Euler’s formulation for density evolution, where the standard continuity equation, Eq. (26), is adopted. For incompressible flow, Euler’s formulation reduces to the condition that the convective velocity field is divergence-free. It is important to note that, for incompressible fluids, the van der Waals and Euler formulations cannot be applied simultaneously, as doing so would lead to an overdetermined system.
@@ IV Dissipation principle
which is consistent with standard theory of diffusion equation. The coefficient $\tau_{0}$ is a positive constant. Substituting Eq. (34) into Eq. (33), integrating by parts, and using the no-flux boundary condition, we obtain
@@
Thus, the coefficient $\tau_{0}$ establishes a connection between the purely dynamic parameter $\mathcal{D}$ (independent of conservative energy) and the energy-dependent term $\frac{\mathrm{d}^{2}e}{\mathrm{d}u^{2}}$ via the relation $\tau_{0}=\mathcal{D}/\frac{\mathrm{d}^{2}e}{\mathrm{d}u^{2}}$.
@@ V Uniform system: Derivation of Euler’s equation and Newton’s law
The free energy density depends on the volume concentration $\phi$ and temperature $T$, and may be expressed as $f=\theta-Ts$, where $\theta$ denotes the internal energy and $s$ represents the entropy.
@@
where the mass diffusivity $D$ is defined in terms of the mobility $M$ as $D=M(\partial^{2}f/\partial\phi^{2})$.
@@ VI.1 Energy formulation and dissipation
This is a standard approach in the literature, except for the inclusion of the pressure term, $\mathcal{P}$. The motivation for introducing $\mathcal{P}$ will be presented below.
The integration over the local free energy density $e(\bm{\phi},\nabla\bm{\phi})$ leads to
$\mathcal{F}(\bm{\phi},\nabla\bm{\phi})=\int_{\Omega}e(\bm{\phi},\nabla\bm{\phi})\mathrm{d}\Omega.$ (53)
By decomposing the entropy $s$ into thermal entropy $s_{T}$, component-mixing entropy $s_{\phi}$, and velocity entropy $s_{u}$, we write the free energy density as
$e=\theta-Ts=\theta-T(s_{T}+s_{\phi}+s_{u})=e_{0}-Ts_{u},$ (54)
where $\theta$ represents the internal energy. The physical meaning of the entropies $s_{\phi}$ and $s_{u}$ is as follows:
  • The component mixing entropy $s_{\phi}$ accounts for the mixing excess volume in the diffusion equation, thereby leading to compressibility.
  • The velocity entropy $s_{u}$ is associated with viscous effects, resulting in velocity dissipation.
The free energy $e_{0}=f-T(s_{T}+s_{\phi})$ does not contain the contribution of the velocity entropy. Here, $f=f_{0}+\sum_{i=1}^{K}\kappa_{i}(\nabla\phi_{i})^{2}$ depicts the van der Waals-Cahn free energy Cahn and Hilliard (1958); Rowlinson (1979) and $\kappa_{i}$ represents the gradient energy coefficient. It should be noted that the entropy increase of $s_{\phi}$ and $s_{u}$ does not necessarily result in heat release or a temperature change of the same magnitude as that associated with thermal entropy. For instance, mass diffusion can lead to an increase in $s_{\phi}$ in an approximately isothermal manner, since the mass diffusivity is much smaller than the thermal diffusivity. Similarly, dissipation of fluid velocity can increase $s_{u}$ nearly isothermally, as the kinematic viscosity of water is one magnitude smaller than its thermal diffusivity. For these physical reasons, we assume isothermal conditions over the time scale of water–mass diffusion in the present work.
@@
$\mathcal{K}=\int_{\Omega}\bm{\zeta}\cdot\mathbf{u}\mathrm{d}\Omega,$ (55)
@@
The $\mathcal{P}=\int_{\Omega}p\mathrm{d}\Omega$ term, often overlooked previously and introduced here, describes the pressure energy. In existing works, the space gradient of pressure is usually employed as the driving force for the fluid motion in the Navier-Stokes equation by using the isothermal Gibbs-Duhem relation
$\mathrm{d}p+\sum_{i=1}^{K}\phi_{i}\mathrm{d}\mu_{i}=0,$ (56)
where $\mu_{i}$ is the chemical potential of species $i$, defined as $\mu_{i}=\delta e/\delta\phi_{i}$. Here, $\delta$ denotes the functional derivative. However, we note that there are two major limitations when using the gradient of pressure $\nabla p$ based on $\mu_{i}=\delta e/\delta\phi_{i}$ and Eq. (56) to derive the momentum evolution.
  • First, the classical Gibbs–Duhem relation, Eq. (56), is derived under a quasi-static assumption, in which the velocity-dependent entropy $s_{u}$ and the associated kinetic energy are not taken into account. Consequently, when $\nabla p$ is used as given in Eq. (56), the role of viscosity in free-energy dissipation or in maximizing the velocity entropy is not clearly defined. One cannot simply introduce a viscous term into the momentum evolution in an ad hoc manner without accounting for its fundamental connection to entropy.
  • Second, the space derivative $\nabla$ and the functional derivative $\delta$ (or $\mathrm{d}$) are not mathematically compatible in most existing works except few references Ridl and Wagner (2018) (see Appendix B).
To solve these two limitations, we derive a modified Gibbs-Duhem relation as follows. The total derivative of the kinetic energy reads, $\delta k=\mathbf{u}\cdot\delta(\rho\mathbf{u})+\rho\mathbf{u}\cdot\delta\mathbf{u}$. The first part $\mathbf{u}\cdot\delta(\rho\mathbf{u})$ is the inertial work which can be absorbed into the first law of thermodynamics in addition to the chemical work $\mu_{i}\delta\phi_{i}$ and mechanical work $p\delta v$. As will be shown in Section VIB, this term, $\mathbf{u}\cdot\delta(\rho\mathbf{u})$ combing with pressure results in the momentum evolution. The second part $\rho\mathbf{u}\cdot\delta\mathbf{u}$ is left into the modified isothermal Gibbs-Duhem relation
$\delta p+\sum_{i=1}^{K}\phi_{i}\delta\mu_{i}+\rho\mathbf{u}\cdot\delta\mathbf{u}=0,$ (57)
as the spirit of Gibbs-Duhem relation is that the total derivative of the energy minus the work associated with the first law of thermodynamics is zero. As $\rho\mathbf{u}\cdot\delta\mathbf{u}=-T\delta s_{u}$ (see Section VIB), we integrate Eq. (57) and obtain a general definition of the pressure as (see Appendix A)
$p=e-\sum_{i=1}^{K}\mu_{i}\phi_{i}+Ts_{u}=e_{0}-\sum_{i=1}^{K}\mu_{i}\phi_{i}.$ (58)
This result provides a clear understanding of the pressure and is fully consistent with previous works Chen and Doolen (1998). It must be noted that the pressure $p$ in Eq. (58) does not include the velocity-dependent entropy, as $e+Ts_{u}$ results in a $s_{u}$ independent energy $e_{0}$.
The $\mathcal{W}$ term in Eq. (52) represents the wall free energy, which is integrated over the wall free energy density $f_{w}$ on the boundary $\Gamma$ as $\mathcal{W}=\int_{\Gamma}f_{w}\mathrm{d}\Gamma$. The wall free energy density $f_{w}$ is responsible for the wetting phenomenon, which has been comprehensively discussed elsewhere Wang and Nestler (2024); Wang et al. (2023); Wang and Nestler (2021).
@@
$\dot{\mathcal{F}}(\bm{\phi},\nabla\bm{\phi})=\int_{\Omega}\sum_{i=1}^{K}\frac{\partial e}{\partial\phi_{i}}\frac{\mathrm{d}\phi_{i}}{\mathrm{d}t}+\frac{\partial e}{\partial\nabla\phi_{i}}\cdot\frac{\mathrm{d}\nabla\phi_{i}}{\mathrm{d}t}\mathrm{d}\Omega.$ (59)
By using the vector calculus $\nabla\cdot(\nabla\phi_{i}\otimes\nabla\phi_{i})=\nabla^{2}\phi_{i}\nabla\phi_{i}+(\nabla\phi_{i}\cdot\nabla)\nabla\phi_{i}$, the definition of materials derivative $\dot{\nabla\phi_{i}}=\partial_{t}\nabla\phi_{i}+\mathbf{u}\cdot\nabla\nabla\phi_{i}$, and integration by parts with no-flux boundary condition (more details can be found in Wang et al. (2023); Zhang et al. (2024)), we obtain the following equations
$\dot{\mathcal{F}}=\int_{\Omega}\sum_{i=1}^{K}\mu_{i}\dot{\phi}_{i}+(\nabla\cdot\underline{\underline{\bm{\Theta}}})\cdot\mathbf{u}\mathrm{d}\Omega;$ (60)
$\mu_{i}=\frac{\delta e}{\delta\phi_{i}}=\frac{\partial e}{\partial\phi_{i}}-\nabla\cdot\frac{\partial e}{\partial\nabla\phi_{i}};$ (61)
@@
The first part of Eq. (60) defines a generalized formulation for the chemical potential, as stated in Eq. (61). At thermodynamic equilibrium, the chemical potential is a constant value throughout the system. In a special case of a zero chemical potential, i.e., $\mu_{i}=0$, Eq. (61) is equivalent to the Euler-Lagrange equation and replicates the Young-Laplace equation; Eq. (61) also provides a physical interpretation for the variational derivative in mathematics. The second part of Eq. (60) defines a stress tensor $\underline{\underline{\bm{\Theta}}}$, which is responsible for the transformation of the potential free energy into the macroscopic kinetic energy Jacqmin (1999). This stress tensor is consistent with the Korteweg stress by evaluating the derivative, $\partial_{\nabla\phi_{i}}e=\kappa_{i}\nabla\phi_{i}$.
@@
where $\mathbf{I}$ is an identity tensor. For the derivation of Eq. (63) and Eq. (64), we have used the chain rule and the definition of the velocity, $\dot{\mathbf{r}}=\mathbf{u}$ without considering relativity effect Landau and Lifshitz (2013). Note that $\mathrm{d}p/\mathrm{d}t=\partial p/\partial t+\mathbf{u}\cdot\nabla p$. For fluids with velocity much less than sound speed, $\partial p/\partial t\approx 0$. Summarizing Eq. (60), Eq. (61), Eq. (62), Eq. (63), and Eq. (64) and collecting all the $\cdot\mathbf{u}$ terms, we obtain the total time derivative of the system energy $\mathcal{E}$ as
$\dot{\mathcal{E}}=\int_{\Omega}\sum_{i=1}^{K}\underbrace{\mu_{i}\dot{\phi_{i}}}_{\text{I}}+\underbrace{[\nabla\cdot(\underline{\underline{\bm{\Theta}}}+p\mathbf{I})+\dot{\bm{\zeta}}]}_{\text{II}}\cdot\mathbf{u}+\underbrace{\bm{\zeta}\cdot\dot{\mathbf{u}}}_{\text{III}}\mathrm{d}\Omega.$ (65)
Note that, at isothermal condition, the total energy is conserved when defining $\Lambda=\mathcal{E}+Ts_{\phi}+Ts_{u}$, where the entropy energies of component and velocity have been added. Within this framework, we write the time derivative of the total energy as
$\frac{\mathrm{d}\Lambda}{\mathrm{d}t}=\int_{\Omega}\sum_{i=1}^{K}\bigg(\mu_{i}\frac{\mathrm{d}\phi_{i}}{\mathrm{d}t}+T\frac{\mathrm{d}s_{\phi_{i}}}{\mathrm{d}t}\bigg)+[\nabla\cdot(\underline{\underline{\bm{\Theta}}}+p\mathbf{I})+\dot{\bm{\zeta}}]\cdot\mathbf{u}+\bigg(\bm{\zeta}\cdot\frac{\mathrm{d}\mathbf{u}}{\mathrm{d}t}+T\frac{\mathrm{d}s_{u}}{\mathrm{d}t}\bigg)\mathrm{d}\Omega=0.$ (66)